OpenStudy (precal):

Calculus - Area between 2 curves y=6x^2-x y=x^2-6x

OpenStudy (precal):

@satellite73

zepdrix (zepdrix):

The frog lady is back! :O oh my

OpenStudy (precal):

ok I have been a little busy. I went back to work......

OpenStudy (precal):

Can you help me please? I think my key is incorrect

zepdrix (zepdrix):

Ah :3

OpenStudy (precal):

my limits according to my key is 0 to 1 but I don't see it

zepdrix (zepdrix):

Looks the the limits are -1 and 0. Your key says 0 and 1..?

OpenStudy (precal):

yes

zepdrix (zepdrix):

So let's find our where they intersect:\[\Large\rm 6x^2-x=x^2-6x\]let's get everything to one side:\[\Large\rm 5x^2+5x=0\]factoring:\[\Large\rm 5x(x+1)=0\]

zepdrix (zepdrix):

Using the Zero Factor Property we get,\[\Large\rm 5x=0 \qquad\implies \qquad x=0\]\[\Large\rm x+1=0\qquad\implies\qquad x=-1\]

OpenStudy (precal):

you are correct, I zoom into the wrong region in my calculator

OpenStudy (precal):

my key is so wrong I was given 4, 25/6, 9/2, 14/3, and 29/6 as choices I was told 25/6 is the solution

zepdrix (zepdrix):

Yah it's a little tricky getting the correct region. https://www.desmos.com/calculator/4b9pmzmtyv There are actually 4 regions right? Only one of them is closed though.

OpenStudy (precal):

I now see my solution should be 5/6

OpenStudy (precal):

I will go with 5/6 and look at my key in the morning. I left my book on my desk in my office.

OpenStudy (precal):

Thanks for your help. :)

zepdrix (zepdrix):

5/6? Ah yes that seems right :) np